OpenStudy (anonymous):

Deep thought granola is 25% nuts and dried fruit. Oat Dreams granola is 10% nuts and dried fruit. How much of deep though and how much oats dream she be mixed to form a 20lbs bath of granola that is 19% nuts and dried fruit?

OpenStudy (anonymous):

set up a table lbs percent product x 25% .25x y 10% .10y _____________________________ 20 19% 20*.19 set up equations x+y = 20 .25x + .1y = 20*.19 solve by substitution

OpenStudy (anonymous):

I am late here....what to sub?

myininaya (myininaya):

if x+y=20, then y=20-x so sub y=20-x whereever you see an x in your second equation

myininaya (myininaya):

wherever you see a y*

OpenStudy (anonymous):

.25*20=5

OpenStudy (dumbcow):

(i) x+y = 20 (ii) .25x + .1y = 20*.19 From (i) y = 20-x replace this with y from (ii) .25x +.1(20-x) = 3.8 distribute .25x +2 -.1x =3.8 add like terms .15x +2 = 3.8 .15x = 1.8 x = 12